40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jun 3 16:46:31 2019, gautam, Update, IOO, IMC not locking PCdrive_RMS.png
    Reply  Tue Jun 4 10:56:31 2019, gautam, Update, IOO, IMC diagnostics IMCdiag.pdf
       Reply  Thu Jun 6 22:11:53 2019, Koji, Update, IOO, IMC diagnostics 
          Reply  Thu Jun 13 18:01:18 2019, aaron, Update, IOO, IMC diagnostics 
             Reply  Fri Jun 14 13:10:00 2019, aaron, Update, IOO, IMC diagnostics 
                Reply  Mon Jun 17 12:37:16 2019, aaron, Update, IOO, IMC diagnostics 
                   Reply  Tue Jun 18 20:35:07 2019, aaron, Update, IOO, IMC diagnostics 
                      Reply  Wed Jun 19 19:12:51 2019, aaron, Update, IOO, IMC diagnostics PMC_Error_Spectrum.pdfPDH_signal.pdfPDH_signal_full.pdf
                         Reply  Mon Jun 24 15:49:05 2019, aaron, Update, IOO, IMC diagnostics sweep_config_updated.png
                            Reply  Tue Jul 2 19:36:18 2019, aaron, Update, IOO, IMC diagnostics PDH_error.pdfPMC_Error_Spectrum.pdf
          Reply  Fri Jun 21 19:36:26 2019, Koji, Update, IOO, IMC diagnostics IMC_error.pdf
             Reply  Sun Jun 23 08:09:53 2019, gautam, Update, IOO, NPRO diagnostics consolidatedOscopeScreenCaps.pdfRINcomp.pdfPLL_OLTF.pdfPLLnoise.pdf
                Reply  Sun Jun 23 09:36:32 2019, gautam, Update, IOO, IMC is locking normally again PMCdiag.pdf
                   Reply  Sun Jun 23 14:43:14 2019, Koji, Update, IOO, IMC is locking normally again 
                   Reply  Mon Jun 24 08:12:10 2019, gautam, Update, IOO, IMC is locking normally again IMCdutycycle.png
                      Reply  Mon Jun 24 11:48:35 2019, gautam, Update, IOO, IMC in-loop error spectra and OLTF IMCfreqNoise.pdfIMC_OLTF.pdfIMC_CMboard.jpg
Message ID: 14693     Entry time: Mon Jun 24 15:49:05 2019     In reply to: 14683     Reply to this: 14721
Author: aaron 
Type: Update 
Category: IOO 
Subject: IMC diagnostics 

aI went to repeat these measurements using the mixer out channel from the servo box, and with a slower sweep for the PDH calibration.

I had trouble getting the PDH signal, here are some notes:

  • I added a 50 Ohm terminator to BNC T on the mixer box. This had been terminated before I started, but I noticed no terminator today.
  • Noticed some distortion of my driving triangle wave if I measured it on ch3 and 4 of the tektronix scope, not present on ch 1/2
  • Initially wasn't finding a signal because I was opening the loop by turning off the Test 1 switch, but this meant the mixer mon on the servo box also did not receive the PDH signal. Instead, I cut the loop with the "BLANK" switch on the PMC screen, which instead blanks out the op amp between the mixer mon and the PMC drive conditioning (so the external drive still reaches the PZT).

attachment 1 is the configuration of the PMC screen when I was trying to get some PDH signal; I did move the DC output adjust to 0V, but found that this led to the output being railed; this makes sense, the op amp at U9 has a negative bias at GND.

Rana came by and gave me some tips.

  • I'd been using the wrong servo board diagram, it should be in D1400221
  • We removed the LP filter from the mixer output (before going to FP1TEST on the servo board), since the board itself already is filtering the IF.
  • We might have observed the thermal locking? See for yourself, the trans and refl signals while sweeping the PZT drive at 5 Hz and 30 Hz respectively are in attachments 2 and 3.
  • Rather than using an SR560, I should use an RF coupler between the mixer and FP1TEST to measure the error signal spectrum. I found a ZFDC-20-5-S+ (0.1-2000 MHz) and sent an SMA cable from the coupled port to channel R of the Agilent 4395.

We finally got the PDH signal again, and I recorded the PDH signal while driving with the following settings on the Siglent function generator.

  • 1.1 Hz triangle wave, 6Vpp, -7Vdc offset, high impedance mode

I tried getting a spectrum using the coupler, the mixer mon is seeing a DC offset though and causing the PZT to rail. Will try to understand why, but in the meantime removing the coupler (still no LP filter) lets us lock the PMC again.

RXA: Kruthi thinks all of our subsequent IMC locking problems are Aaron's fault (she was quick to give him up as soon as the thumb screws were tightened...)

Attachment 1: sweep_config_updated.png  28 kB  Uploaded Mon Jun 24 18:25:16 2019  | Hide | Hide all
sweep_config_updated.png
ELOG V3.1.3-